### v1

# StrandHead: Text to Strand-Disentangled 3D Head Avatars Using Hair Geometric Priors

Abstract
While haircut indicates distinct personality, existing avatar generation methods fail to model practical hair due to the general or entangled representation. We propose StrandHead, a novel text to 3D head avatar generation method capable of generating disentangled 3D hair with strand representation. Without using 3D data for supervision, we demonstrate that realistic hair strands can be generated from prompts by distilling 2D generative diffusion models. To this end, we propose a series of reliable priors on shape initialization, geometric primitives, and statistical haircut features, leading to a stable optimization and text-aligned performance. Extensive experiments show that StrandHead achieves the state-of-the-art reality and diversity of generated 3D head and hair. The generated 3D hair can also be easily implemented in the Unreal Engine for physical simulation and other applications.
Figure 1: We propose StrandHead, a text-driven framework for generating strand-disentangled 3D head avatars that feature high-fidelity facial details and strand-based hair. By accurately capturing the internal geometry of hair strands, our approach seamlessly supports flexible hairstyle transfer and editing, as well as physics-based rendering and simulation.
1 Introduction
Creating 3D head avatars is crucial for many applications including digital telepresence, gaming, movies, and AR/VR. Traditional approaches heavily depend on manual efforts which is time-consuming and labor-intensive. Recent advances [55, 84, 93, 79] achieve automatic 3D head generation with the learning-based paradigm. Although these methods provide reliable 3D head geometry, they require high-cost 3D training data, which significantly constrains their generalization capabilities and limits the potential applications.
With the rapid development of text-driven generation methods [57, 50], creating 3D head avatars from given prompts becomes possible. Recent studies [1, 12, 99, 89] build upon the standard text-to-3D generation frameworks, achieving low-cost and high-fidelity performance by integrating with 3D parametric head models [4, 26, 82]. While focusing on facial geometry and texture modeling, these methods utilize holistic meshes or NeRF [39] to represent 3D haircut, failing to capture the internal geometric structure of hair strands, i.e., 3D curves. This limitation not only significantly reduces the realism of the generated avatars but also makes them incompatible with existing strand-based rendering and simulation systems [10, 2, 9].
To model 3D haircuts, recent efforts achieve strand-level hairstyle reconstruction [58, 66, 94] or generation [97, 13] using VAEs or parametric models. Despite their impressive performance, they require constrained multi-view images or manually latent space searching, failing to freely create 3D hair under semantics of textual prompts. The most recent work HAAR [67] pioneers text-to-strand generation by training a text-conditioned hair map diffusion model. However, due to the high cost and limited scale of available 3D hair data, HAAR struggles to generate hairstyles outside of the training set. In addition, HAAR does not model hair texture or geometry that adapts to a specific generated human head, which limits its applicability in downstream tasks.
In this paper, we propose StrandHead, a novel framework that generates strand-disentangled 3D head avatars from text descriptions. As illustrated in Fig. 1, the generated head avatars exhibit detailed 3D faces with high-fidelity hair strands and diverse 3D haircuts, enabling seamless and flexible haircut transfer, editing, and physical-based simulation. Without requiring any hair training data, we create 3D hair by distilling the 2D generative diffusion models. This is achieved by employing a series of key insights on hair geometric priors. First, inspired by the cylindrical structure of hair strands [19, 35], we propose a novel differentiable prismatization algorithm. This algorithm facilitates the efficient, flexible, and differentiable transformation of strands into watertight prismatic meshes with customizable thickness and defined lateral edges, enabling the use of mesh-based renderers, loss functions, and models for detailed strand-level modeling tasks. Second, leveraging the distribution patterns of 3D hair orientation and curvature, we propose reliable consistency regularization losses for robust generation. These losses enhance the cosine similarity between neighboring strand orientations and constrain the global hair curvature. Extensive experiments demonstrate that StrandHead outperforms the state-of-the-art (SOTA) methods in both head and hair generation tasks, and supports flexible haircut transfer and editing, as well as physical-based rendering and simulation.
Our main contributions are summarized as follows:
  • We propose StrandHead, a novel framework for generating high-fidelity 3D head avatars with disentangled strand-based hair. To the best of our knowledge, StrandHead is the first work to generate strand-disentangled 3D head from pre-trained 2D diffusion models.
  • We propose a differentiable prismatization algorithm that converts hair strands into watertight prismatic meshes, facilitating the use of mesh-based renderers, losses, algorithms, and models in strand-level deep learning tasks.
  • Inspired by statistical hair features, we propose orientation consistency and curvature regularization losses to supervise both local and global hair shape, enabling reliable and realistic hairstyle generation.
2 Related Work
Text-to-3D General Object Generation. Inspired by the success of text-to-image (T2I) generation [43, 53, 54, 59, 57, 91], many studies have explored using pre-trained vision-language models [52, 57] to achieve text-guided zero-shot 3D content generation. Early methods [40, 38, 61, 18, 71] employ the CLIP model [52] to supervise the alignment of 3D representations with prompts. DreamFusion [50] introduces the Score Distillation Sampling (SDS) loss, significantly enhancing the fidelity of generated 3D content by leveraging more powerful pre-trained diffusion models [57]. Subsequent works further advance text-driven 3D generation by improving 3D representations [5, 6, 85], optimization strategies [29, 70, 86], SDS loss [75, 28], and diffusion models [27, 51, 65]. Despite these advancements, current approaches focused on general content generation do not fully leverage the extensive prior knowledge of human heads and hair, limiting their ability to generate realistic, high-quality 3D head avatars with strand-based hair.
Text-to-3D Head Avatar Generation. CLIPFace [1], T2P [92] and Describe3D [78] pioneer zero-shot text-driven 3D head generation by combining CLIP [52] with 3D parametric head models [4, 26, 8, 82]. With the introduction of SDS loss [50], DreamFace [90] and FaceG2E [80] leverage pre-trained T2I diffusion models to greatly improve generation fidelity within specific domains. HeadEvolver [72] enhances the expressiveness of head mesh deformations by introducing vector fields, while HumanNorm [17], HeadSculpt [12], and HeadArtist [30] employ DMTet [63] instead of traditional meshes to capture geometric details. HeadStudio [99] incorporates 3DGS [20] to produce realistic and animatable 3D heads. Despite these advancements, these approaches treat the head and hair as a holistic model, limiting support for downstream applications such as hairstyle transfer and editing. To enable flexible head-hair-disentangled generation, TECA [89] independently represents the head with mesh and the hair with NeRF [39]. However, these methods focus only on the realistic external appearance of hair without modeling its internal geometric structure, restricting their application for strand-level physics-based rendering and simulation.
Strand-Based Hair Creation. Automatic strand-based hair modeling has attracted substantial attention from both industry [10, 2, 9] and academia [87, 34, 3, 14], as it enables downstream physics-based rendering and simulation Such methods can be roughly divided into two categories: (1) reconstructing hair from input images or videos and (2) generating hair based on control conditions. Works on the former one [15, 96, 60, 42, 21, 58, 76, 83, 64, 66, 94, 33, 88, 77, 69, 95, 98, 36] relies on classical 3D reconstruction frameworks [62, 46, 37, 39, 20] and incorporates hair-specific features [45] to realize accurate 3D strand reconstruction. However, these methods depend on constrained multi-view images, making hair creation costly and challenging in less controlled settings. Recently, 3D hair generation approaches [97, 13] have emerged using GANs or parametric model, but they struggle with controllable generation. The most relevant work to ours is HAAR [67], which pioneers text-guided strand generation using diffusion models. However, HAAR relies on high-cost 3D hair data which is limited in diversity, restricting its ability to generate novel haircut beyond the training data.
Compared to previous methods, StrandHead generate 3D heads with fine geometry and lifelike textures, as well as realistic textured strand-based hairstyles, and the entire framework does not require any hair training data. We summarize the main differences between our approach and related approaches in Tab. 1.
Task Method Head-Hair- Decoupled Strand-Based Hair Geometry & Texture No Training Data
Text-to-Head [12, 30, 99, 17] [89]
[…]
Table 1: Comparison with current related methods.
Figure 2: Strandhead consists of two stages: (a) Under the constraints of the human-specific diffusion model and the FLAME-volving prior loss, StrandHead generates a detailed and reasonable bald head. (b) By introducing a differentiable prismatization algorithm, orientation consistency loss and curvature regularization loss inspired by hair geometric priors, StrandHead achieves diverse and realistic strand-accurate hair creation without any requiring hair training data.
3 Method
3.1 Preliminaries
FLAME [26] is a 3D parametric head model that represents the human head shape, pose, and expression using a compact set of parameters. Given a set of shape parameters $\beta$, pose parameters ${\theta}_{\text{head}}$, and expression parameters $\psi_{\text{exp}}$, FLAME can create a 3D head mesh ${M}$.
Score Distillation Sampling (SDS) [50] uses a pre-trained T2I model to guide the optimization of a 3D representation based on the input text. For an input text prompt $y$, a 3D representation parameterized by $\theta$, and a T2I diffusion model with parameters $\phi$, the SDS loss is defined as follows:
$\nabla_{\theta}\mathcal{L}_{SDS}=\mathbb{E}_{t,\epsilon}\left[w(t)({\epsilon}_{\phi}(x_{t};y,t)-\epsilon)\frac{\partial x}{\partial\theta}\right],\vspace{-0.3cm}$ (1)
where $x=g(\theta)$ is the image rendered from $\theta$ by a differentiable renderer $g$, $t$ is the time step, $x_{t}=x+\epsilon$ is a noised version of $x$, ${\epsilon}_{\phi}(x_{t};y,t)$ is the denoised image, and $w(t)$ is a weighting function. For simplicity, we omit $w(t)$ in the following formulas.
DMTet [63] is a hybrid representation that marrys a signed distance function (SDF) and a differentiable Marching Tetrahedral layer to convert implicit SDF to an explicit mesh. Thanks to multi-resolution hash encoding [41], the SDF can be implemented in a lightweight manner.
Neural Scalp Textures (NST) [58] is an efficient hair representation where each pixel stores a feature vector conveying the shape information of a single strand at the corresponding scalp location. With the help of a pre-trained hair strand generator [58] $G$, the low-dimensional 2D neural scalp texture $T$ can be decoded into high-dimensional, dense 3D strand polylines $S=\{s^{i}\}_{i=1}^{N_{s}}$, where $N_{s}$ is the number of strand polylines, and $s^{i}=\{p^{i}_{j}\}_{j=1}^{N_{p}}$ is a polyline formed by $N_{p}$ 3D points. This process can be formulated as follows: $S=G(T)$.
3.2 Overview
Given a text prompt, StrandHead aims to create a high-quality 3D head avatar with strand-disentangled textured hair. Fig. 2 provides an overview of our pipeline. We first produce a high-fidelity and reasonable 3D bald head (Sec. 3.3). Then, we generate realistic, reasonable strand-based hair with a lifelike appearance by introducing a differentiable prismatization algorithm, two prior-driven losses and strand-aware texture field (Sec. 3.4).
3.3 Bald Head Generation
Head Geometry Modeling. To sculpt detailed geometry, we follow HeadArtist [30] and HeadSculpt [12] using DMTet [63] as our 3D representation and initialize it with the FLAME mesh $M_{\text{init}}$. After initialization, we leverage Nvdiffrastt [22] and SDS loss to model the geometry and texture of the human head in a decoupled manner. We employ a series of human-specialized diffusion models proposed by HumanNorm [17] to further improve the fidelity of the generated head. Specifically, we optimize the head DMTet parameterized by ${\theta}_{\text{h}}$ using the following losses:
$\nabla_{{\theta}_{\text{h}}}\mathcal{L}^{\text{hn}}_{SDS}$ $=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{\text{hn}}}(n^{\text{h}}_{t};y_{\text{h}},t)-\epsilon)\frac{\partial n^{\text{h}}}{\partial{\theta}_{\text{h}}}\right],$ (2)
$\nabla_{{\theta}_{\text{h}}}\mathcal{L}^{\text{hd}}_{SDS}$ $=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{\text{hd}}}(d^{\text{h}}_{t};y_{\text{h}},t)-\epsilon)\frac{\partial d^{\text{h}}}{\partial{\theta}_{\text{h}}}\right],\vspace{-0.5cm}$ (3)
where $y_{\text{h}}$ is the input bald head text, ${\phi}_{\text{hn}}$ and ${\phi}_{\text{hd}}$ are the human-specialized normal-adapted and depth-adapted diffusion models, respectively, and $n^{\text{h}}$ and $d^{\text{h}}$ are the rendered normal and depth maps of the bald head.
Additionally, inspired by Barbie [68], we impose a FLAME-evolving prior loss to prevent unnatural geometry (e.g., overly large ears or overly pointed head tops). In particular, we freeze the FLAME shape parameters $\beta$ but improve $M_{\text{init}}$ to evolvable $\hat{M}_{\text{init}}$ by adding learnable vertex-wise offsets, and periodically fit $\hat{M}_{\text{init}}$ to the head DMTet every $\delta$ iterations, enabling the learned offsets to model complex facial features (e.g., beards and wrinkles, see the Supp. Mat. for loss details). During evolution, $\hat{M}_{\text{init}}$ gradually captures rich geometric details, and provide a reliable and diverse head prior for geometry generation using the following formula:
$\mathcal{L}_{\text{prior}}=\sum_{p\in P}\left\|s_{\theta_{\text{h}}}(p)-s_{\hat{M}_{\text{init}}}(p)\right\|_{2}^{2},\vspace{-0.4cm}$ (4)
where $s_{\theta_{\text{h}}}$ and $s_{\hat{M}_{\text{init}}}$ are the SDF functions of the head DMTet and $\hat{M}_{\text{init}}$, and $P$ is a set of randomly sampled points. In summary, the loss function for optimizing the head geometry is as follows:
$\mathcal{L}_{\text{head-geo}}=\mathcal{L}^{\text{hn}}_{SDS}+\mathcal{L}^{\text{hd}}_{SDS}+\lambda_{\text{prior}}\mathcal{L}_{\text{prior}}.\vspace{-0.2cm}$ (5)
Head Texture Modeling. Given the generated head geometry generated, we fix it and utilize a texture field ${\psi}_{\text{h}}$ which maps a query position to its color to generate head appearance. Specifically, we construct this field using a MLP network with multi-resolution hash encoding [41], and we optimize it with the following loss function:
$\nabla_{{\psi}_{\text{h}}}\mathcal{L}^{\text{hc}}_{SDS}=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{\text{hc}}}(c^{\text{h}}_{t};n^{\text{h}},y_{\text{h}},t)-\epsilon)\frac{\partial c^{\text{h}}}{\partial{\psi}_{\text{h}}}\right],\vspace{-0.2cm}$ (6)
where ${\phi}_{\text{hc}}$ is a human-specialized normal-conditioned diffusion model, and $c^{\text{h}}$ represents the rendered color image of the generated head. Since the vanilla SDS loss often leads to color oversaturation, we replace it with the following MSDS loss [17] to further enhance the texture’s realism in later iterations of texture optimization:
$\nabla_{{\psi}_{\text{h}}}\mathcal{L}^{\text{hc}}_{MSDS}=\mathbb{E}_{t,\epsilon}\left[(H(c^{\text{h}}_{t};n^{\text{h}},y_{\text{h}},t)-\epsilon)\frac{\partial c^{\text{h}}}{\partial{\psi}_{\text{h}}}\right],\vspace{-0.2cm}$ (7)
where $H(\cdot)$ denotes the multi-step operation function.
3.4 Strand-Based Hair Generation
Hair Initialization. To achieve accurate and diverse hair initialization, we utilize ChatGPT [44] to select the 20 most representative hairstyles from the USC-HairSalon Dataset [15]. We then use the following loss to optimize NST to fit the selected hairstyle:
${\mathcal{L}}_{\text{fit}}=\sum^{N_{s}}_{i=1}\sum^{N_{p}}_{j=1}\|\hat{p}^{i}_{j}-p^{i}_{j}\|_{2}+\lambda_{\text{ori}}(1-\hat{o}^{i}_{j}\cdot o^{i}_{j})+\lambda_{\text{cur}}\|\hat{c}^{i}_{j}-c^{i}_{j}\|_{1},$ (8)
where $\hat{p}^{i}_{j}$ and ${p}^{i}_{j}$ are the position of the $j$-th point on the $i$-th polyline of the GT and generated hair, respectively, and $\hat{o}^{i}_{j}$, $o^{i}_{j}$, $\hat{c}^{i}_{j}$, and $c^{i}_{j}$ denote their orientation and curvature, respectively. Here, orientation $o^{i}_{j}=(p^{i}_{j+1}-p^{i}_{j})/\|p^{i}_{j+1}-p^{i}_{j}\|_{2}$ represents the direction of change in strand position, and curvature $c^{i}_{j}=\|o^{i}_{j}-o^{i}_{j-1}\|_{2}$ denote the rate of change in strand orientation. Given the hair description, we select an optimal pre-trained NST as initialization and further optimize it in the following stages.
Differentiable Prismatization. Actually, as there is no reliable strand-based differentiable renderer, performing end-to-end hair strands optimization becomes challenging. Further, since we have no constrained multi-view images, we cannot directly model the hair geometry and appearance using 3DGS [20] or line-based renderer [58], as existing hair reconstruction methods [58, 69, 33, 88]. A feasible alternative is to first convert the strands into quad meshes [87], similar to NeuralHaircut [66], and then leverage mesh optimization frameworks to model hair strands. However, this non-watertight stripe-like mesh easily produces ambiguous normal maps or excessively thin sides, which significantly reduce the stability of optimization (Sec. 4.4). Therefore, a more stable strand-to-mesh conversion method with superior optimization properties is urgently necessary.
To this end, we propose a novel differentiable prismatization algorithm inspired by the internal structure of hair (i.e., hair fiber is a dielectric cylinder covered with tilted scales and with a pigmented interior [19, 35]). This algorithm can efficiently and differentiably convert strands into watertight prismatic meshes with arbitrary thickness and lateral edges to approximate the cylindrical structure of hair, enabling the use of mesh-based renderers and algorithms for strand-level modeling. Specifically, given a hair strand $s$, our differentiable prismization algorithm converts it into a watertight prismatic mesh with $K$ lateral edges and radius $R$ through the following five steps:
  1. 1.
    Compute the Initial Normal Vector: Determine a normal to the hair strand $s$ by taking the cross product of its orientation with a non-collinear reference point (typically the center of the head).
  2. 2.
    Generate $K$ Rotated Normals: Rotate this normal around the axis defined by the strand’s orientation $K$ times, each by $\frac{360^{\circ}}{K}$, to produce $K$ normals.
  3. 3.
    Translate to Form Lateral Edges: Translate $s$ along each of these $K$ normals by $R$, generating $K$ lateral edges of the prism.
  4. 4.
    Construct Lateral Faces: Connect the adjacent lateral edges’vertices to form the $K$ lateral faces of the prism.
  5. 5.
    Construct Top and Bottom Faces: Connect the vertices at the ends of the lateral edges to form the top and bottom faces of the prism, completing the conversion from a hair strand to a watertight prismatic mesh.
Figure 3 shows an example of converting a hair strand into an octagonal prism mesh. Importantly, the proposed differentiable prismatization algorithm can be easily implemented on GPU, achieving flexible and fast prismatization of hair strands and paving a new way for hair modeling. The algorithm details are shown in the Supp. Mat.
Figure 3: Illustration of the process of converting a hair strand into an octagonal prism mesh using the differentiable prismatization algorithm.
Hair Geometry and Texture Modeling. With the help of our proposed differentiable prismatization algorithm, we create the hair geometry and texture in two stages as Sec. 3.3 using the following losses:
$\nabla_{T}\mathcal{L}^{\text{hn}}_{SDS}$ $=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{\text{hn}}}(n^{\text{h+s}}_{t};y_{\text{h}},t)-\epsilon)\frac{\partial n^{\text{h+s}}}{\partial T}\right],$ (9)
$\nabla_{T}\mathcal{L}^{\text{hd}}_{SDS}$ $=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{\text{hd}}}(d^{\text{h+s}}_{t};y_{\text{h}},t)-\epsilon)\frac{\partial d^{\text{h+s}}}{\partial T}\right],$ (10)
$\nabla_{{\psi}_{\text{s}}}\mathcal{L}^{\text{hc}}_{SDS}$ $=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{\text{hc}}}(c^{\text{h+s}}_{t};n^{\text{h+s}},y_{\text{h}},t)-\epsilon)\frac{\partial c^{\text{h+s}}}{\partial{\psi}_{\text{s}}}\right],$ (11)
$\nabla_{{\psi}_{\text{s}}}\mathcal{L}^{\text{hc}}_{MSDS}$ $=\mathbb{E}_{t,\epsilon}\left[(H(c^{\text{h+s}}_{t};n^{\text{h+s}},y_{\text{h}},t)-\epsilon)\frac{\partial c^{\text{h+s}}}{\partial{\psi}_{\text{s}}}\right],\vspace{-1cm}$ (12)
where $n^{\text{h+s}}$, $d^{\text{h+s}}$, and $c^{\text{h+s}}$ are the rendered normal, depth, and color maps of the head $\theta_{\text{h}}$ with strand-based hair prismatic mesh $M_{p}=DP(G(T))$, respectively. Here $DP(\cdot)$ denotes the differentiable prismatization operation function, ${\psi}_{\text{s}}$ is the strand texture field.
Figure 4: The distribution of $O_{\text{ori}}$ and $C_{\text{mean}}$ in the USC-HairSalon dataset [15]. The results indicate that (1) neighboring strand orientations are highly consistent; (2) strand curvature is strongly and positively related to the haircut curliness.
Prior-Driven Losses. However, due to the excessive flexibility of hair strands, relying solely on the SDS loss leads to unnatural strand orientations and unrealistic hairstyles (Sec. 4.4). Further, the lack of constrained multi-view images prevents us from leveraging hair-specific features [45] to provide powerful supervision as in previous works [58].
Figure 5: Examples of high-fidelity and diverse 3D heads and strand-accurate haircuts generated by our method. The upper visualization includes rendered color and normal maps of the head and hair prism meshes. The lower visualization shows the physics-based rendering result using Blender [10]. Please zoom in for detailed views, and refer to the Supp. Mat. for video demonstrations.
To address this issue, we propose two straightforward yet solid prior-driven losses based on observations of the hair strand orientation and curvature to ensure the rationality of the generated hair strands. Specifically, we observe two geometric properties of hair: (1) neighboring strand orientations are highly consistent, and (2) strand curvature is strongly and positively correlated with the curliness of the hairstyle. To validate the properties, we calculate the cosine similarity of adjacent hair strand orientations $CS_{\text{ori}}$ and the average curvature $C_{\text{mean}}$ of 343 hairstyles in the USC-HairSalon dataset [15] using the following equations and visualize their distributions and some examples in Fig. 4:
$CS_{\text{ori}}$ $=\frac{1}{N_{s}N_{p}}\sum_{i=1}^{N_{s}}\sum_{j=1}^{N_{p}}\sum_{k\in A(i)}\frac{o^{i}_{j}\cdot o^{k}_{j}}{|A(i)|},$ (13)
$C_{\text{mean}}$ $=\frac{1}{N_{s}N_{p}}\sum^{N_{s}}_{i=1}\sum^{N_{p}}_{j=1}c^{i}_{j},$ (14)
where $A(i)$ is the set of adjacent strands for the $i$-th strand, and $|A(i)|$ is the number of neighboring strands.
As illustrated, over 95% of hairstyles exhibit an orientation similarity above 0.9, and the average curvature is significantly positively related to curliness, which successfully validates the properties. Based on these observations, we introduce orientation consistency loss and curvature regularization to guide the local and global strand shapes, formulated as follows:
$\mathcal{L}_{\text{ori}}$ $=1-CS_{\text{ori}},$ (15)
$\mathcal{L}_{\text{cur}}$ $=\|C_{\text{mean}}-C_{\text{target}}\|_{1},$ (16)
where $C_{\text{target}}$ represents the target average curvature set according to the input hair description. These two losses ensure the rationality of generated strands by supervising the consistency of orientations between adjacent strands and the overall hairstyle’s curvature.
Additionally, we introduce a series of geometry-aware losses $\mathcal{L}_{bbox}$, $\mathcal{L}_{face}$, and $\mathcal{L}_{colli}$, which prevent the hair from exceeding the bounding box, obscuring the face, and colliding with the head, respectively. Details are provided in the Supp. Mat. The final loss function for optimizing hair geometry is expressed as follows:
$\mathcal{L}_{\text{hair-geo}}=$ $\mathcal{L}^{\text{hd}}_{SDS}+\lambda^{\text{hn}}_{SDS}\mathcal{L}^{\text{hn}}_{SDS}+\lambda_{\text{ori}}\mathcal{L}_{\text{ori}}+\lambda_{\text{cur}}\mathcal{L}_{\text{cur}}$
$+\lambda_{\text{bbox}}\mathcal{L}_{\text{bbox}}+\lambda_{\text{face}}\mathcal{L}_{\text{face}}+\lambda_{\text{colli}}\mathcal{L}_{\text{colli}}.$ (17)
Strand-Aware Texture Field. Since the complexity of hair textures, a basic texture field cannot fully capture the lifelike appearance of strands (Sec. 4.4). Therefore, we propose a strand-aware texture field that better models high-frequency color variations. Specifically, we introduce two improvements to the basic texture field: first, we replace Euclidean coordinates with scalp UV coordinates which are more uniformly distributed. Second, we incorporate strand orientations as additional input information to model orientation-dependent texture. These enhancements enhance the realism of the generated results by better creating the high-frequency appearance variations, but also ensure consistent colors across different faces of a single prismatic mesh since the input features are strand-based. For more details, please refer to the Supp. Mat.
3.5 Implementation Details
Our approach utilizes PyTorch [47] and ThreeStudio [11] for implementation. Experiments are carried out on an Ubuntu server with A6000 GPUs. Starting with a textual description of a head with hair, such as “A man with brown hair”. we initially create a bald head using the prompt “A bald man” and then generate hair based on “A man with brown hair”. Further implementation details are available in the Supp. Mat.
Figure 6: Qualitative comparisons with the SOTA methods. Since TECA [89] uses the vanilla NeRF to represent hair, rendering normals is not supported. HAAR [67] generates only the geometry of hair strands, so we first convert the strands into prismatic meshes using differentiable prismatization and then utilize TEXTure [56] to generate texture for visualization and comparison.
Figure 7: Qualitative comparison with HAAR [67]. For better visual comparison, we interpolate the hairstyles to approximately 10,000 strands and apply a consistent appearance. Since HAAR does not model heads, its generated results frequently display hair-head collisions, highlighted within the black box.
4 Experiments
Some examples of 3D head models with strand-based hair generated by StrandHead are shown in Fig. 5. Due to space limitations, only the key experimental settings and results are presented here. For more results and details, please refer to our Supp. Mat.
4.1 Experimental Settings
Baselines. We compare StrandHead with the SOTA methods for head avatar and haircut generation. The baseline methods for head generation include text-to-holistic-head works (HeadArtist [30], HumanNorm [17], and HeadStudio [99]) and text-to-decoupled-head work (TECA [89]). For text-to-hair generation, we consider general text-to-3D methods (MVDream [65], GaussianDreamer [85], LucidDreamer [28], and RichDreamer [51]), strand-specific text-to-hair methods (HAAR [67]), and hair-disentangled text-to-head methods (TECA [89]).
Dataset Construction. To create the dataset, we use ChatGPT[44] to randomly generate 30 text descriptions of heads with hair. These include 30 descriptions for assessing head generation and 30 for evaluating hair generation. The full list of prompts can be found in the Supp. Mat.
Evaluation Metrics. Current text-to-3D methods typically use CLIP-based metrics to assess text-image alignment and the quality of generated outputs. However, research has shown [16, 32] that CLIP-based metrics fall short in capturing the fine-grained alignment between 3D content and input prompts, as further validated by our experiments in the Supp. Mat. To address this limitation, and drawing inspiration from Progressive3D [7] and Barbie [68], we use BLIP-VQA [24, 23] and BLIP2-VQA [25, 23] to evaluate both current methods and StrandHead’s generative capabilities. Specifically, we convert each prompt into two questions to separately verify head and hair. We then feed the rendered image of the generated 3D content into the VQA model, asking each question in sequence and using the probability of a “yes” response as our evaluation metric. For example, the head prompt “A man with black hair” is transformed into “Is the person in the picture a man?” and “Does the person in the picture have black hair?” The hair prompt “Black hair” becomes “Is the object in the picture black hair?” Finally, we conduct a user study by randomly selecting 10 generated examples and asking 30 volunteers to assess (1) generation quality and (2) text-image alignment, and select the best methods.
Method BLIP-VQA $\uparrow$ BLIP2-VQA $\uparrow$ GQP $\uparrow$ TAP $\uparrow$
HeadArtist 0.7667 0.9667 1.00 2.33
HeadStudio 0.7833 0.8833 3.33 3.67
HumanNorm 0.7000 0.9500 7.67 7.67
TECA 0.7333 0.9500 34.33 28.33
StrandHead (Ours) 0.8500 0.9667 53.67 58.00
MVDream 0.9000 0.8333 24.67 20.00
GaussianDreamer 0.3333 0.3667 5.33 3.00
LucidDreamer 0.8000 0.9333 5.33 5.00
RichDreamer 0.8333 0.7667 4.00 5.67
TECA 0.7000 0.7667 1.67 3.67
HAAR 0.6333 0.2000 1.33 2.33
StrandHead (Ours) 0.9000 0.9000 57.67 60.33
Table 2: Quantitative comparisons with the SOTA methods. The best and second-best results are highlighted in bolded and underlined, respectively. GQP: generation quality preference (%), TAP: text-image alignment preference (%).
4.2 Comparisons of Head Generation
As illustrated in the upper section of Tab. 2, our approach outperforms all comparison methods across every evaluation metric. The qualitative results on the left side of Fig. 6 further highlight the advantages of our method. Compared to other head generation methods, the head avatars generated by StrandHead exhibit not only more refined facial geometry and texture details but separable strand-accurate hair with a plausible appearance that integrates seamlessly with physics-based rendering and simulation systems. To the best of our knowledge, StrandHead is the first head generation framework to achieve strand-level hair modeling, a capability that holds substantial potential for advancing human-centric 3D AIGC applications in the industry.
4.3 Comparisons of Hair Generation
As shown in the lower part of Tab. 2, our approach surpasses other methods across most evaluation metrics, and is only slightly lower than LucidDreamer (the latest open-source text-to-3D method) on BLIP2-VQA, ranking the second. The qualitative comparisons on the right side of Fig. 6 display sample generation results. Compared to general text-to-3D methods, our approach more precisely captures the internal structure of hair strands, resulting in realistic haircut geometry and appearance without incorporating incorrect content such as parts of the human head.
Unlike the text-to-strand approach HAAR [67], our approach does not require 3D hair training data, enabling it to better generate uncommon hairstyles (e.g., slicked-back and side-swept hairstyles), as further validated by the physics-based rendering comparison in Fig. 7. Additionally, StrandHead can generate both hair textures and head models, which not only extends the range of downstream applications but also avoids hair-head collisions (see Fig. 7).
Figure 8: Ablation study on (a) strand-level optimization, (b) orientation consistency loss, (c) curvature regularization loss, (d) differentiable prismatization, and (e) strand-aware texture field.
4.4 Ablation Study
Effect of Strand-Level Optimization. As illustrated in Fig. 8-(a), under dual constraints of SDS loss and prior-driven losses, the optimized hair better aligns with the text and integrates harmoniously with the human head.
Effect of Orientation Consistency Loss. Fig. 8-(b) demonstrates that incorporating $\mathcal{L}_{\text{cur}}$ provides strong guidance for generating realistic and natural hair.
Effect of Curvature Regularization Loss. As shown in Fig. 8-(c), our $\mathcal{L}_{\text{cur}}$ preserves the desired curliness by constraining the strand curvature.
Effect of Differentiable Prismatization. As shown in Fig. 8-(d), compared to the quad mesh used by NeuralHaircut [66] our proposed differentiable prismatization algorithm enables more stable gradient backpropagation for reliable strand-based hair optimization, while also avoiding abnormal normal issues caused by non-watertight meshes.
Effect of Strand-Aware Texture Field. As shown in Fig. 8-(e), our proposed strand-aware texture field accurately models high-frequency appearance details by switching coordinate spaces and incorporating orientation information, resulting in more realistic strand textures.
Figure 9: Applications of StrandHead.
4.5 Applications
As illustrated in Fig. 9-(a), StrandHead enables the flexible combination of any head with any hairstyle, greatly enhancing the playability and reusability of head generation. Furthermore, due to our precise strand-level hair modeling, we can not only edit hairstyles by manipulating strand geometry and texture but also leverage mature industrial software to achieve physics-based realistic rendering and simulation, as shown in Fig. 9-(b) and (c).
5 Conclusion
In this paper, a novel framework StrandHead is proposed to generate realistic hair-disentangled 3D head avatar using text prompts. StrandHead generates disentangled strand-based hair only from the pre-trained 2D generative diffusion models, requiring no extra 3D hair data for supervision. StrandHead achieves this goal mainly using two approaches: (1) proposing a novel differentiable prismatization method that transforms the hair strands into meshes of octagonal prisms, and (2) using the prior-driven losses inspired by the statistical features of 3D hair data. The former one produces geometry more similar to the original hair strands, and makes the optimization available to mesh-based renderers with significantly less ambiguity. Besides, the latter one well constrains the optimization with reliable guidance to generate reasonable 3D haircut. Extensive experiments demonstrate that StrandHead obtains the state-of-the-art performance on the generated 3D head avatars and haircuts. The generated results are easily implemented in industrial softwares to produce physical simulation, and high-fidelity rendering.